When the birthday cake was about to served,you were told that you could have 0.6,3/5,60%6%.which 3will give you the same size portion


Answer by thejackal(72)

0.6,3/5,60%
because
0.6 = 6/10 which when expressed as a percentage = (6/10)*100 = 60%
3/5 = 0.6 = above
and 60% which is self explanatory